1. You are a quantitative geneticist campaigning to raise awareness for proper genetic counselling about sickle cell disease in Saudi Arabia. You conduct a study to determine the frequency of sickle cell genotype (ss) among the citizens living in Khobar, Saudi Arabia. You randomly sample and screen 500 subjects for the sickle cell genotype for the age group of 1-50 years and find that 369 of those individuals are homozygous dominant (SS), 115 are heterozygous (Ss), and 16 are homozygous recessive (ss). 



  1. Calculate the allele frequencies (show your work). Hint, use:  2 x (SS) + (Ss)/2 x 500 to calculate your (S) allele frequency. Subtract that result from 1 to get your (s) allele frequency. 
1
Expert's answer
2021-03-16T04:56:05-0400

The frequency of homozygous dominant individuals equals:

f(SS) = n(SS) / n(total) = 369 / 500 = 0.738

The frequency of homozygous recessive individuals equals:

f(SS) = n(SS) / n(total) = 16 / 500 = 0.032

From here, the frequency of heterozygous individuals equals:

f(2Ss) = 1 - 0.738 - 0.032 = 0.23

As a result:

2Ss = 0.23 and S + s = 1. From here:

2S(1-S) = 0.23

S2 - S + 0.115 = 0

S = 0.87

s = 1 - 0.87 = 0.13


Answer: 0.87 and 0.13